OpenStudy (anonymous):

What is 10% more than 90

OpenStudy (anonymous):

90+.10(90) 90+9 =99

OpenStudy (anonymous):

my answer is not right.

OpenStudy (skullpatrol):

10% is just another way to write 10(0.01) since the % symbol means "divided by 100"

OpenStudy (skullpatrol):

10(0.01) = 0.1

OpenStudy (anonymous):

oh i am sorry.

OpenStudy (skullpatrol):

So 10% of 90 is 0.1*90 = 9

OpenStudy (skullpatrol):

What is 9 more than 90? 9 + 90 = 99

OpenStudy (anonymous):

10% of 90 is 9, therefore: 90 + 9 = 99 Your answer is 99 :